# Constant Rate Scheduling: Constant-Rate Distributional Change for Efficient Training and Sampling in Diffusion Models

### v3

@@ 4.2 Constant Rate Scheduling
$D(t,t+\Delta t)\simeq\left.\frac{\partial D(t,t^{\prime})}{\partial t^{\prime}}\right|_{t^{\prime}=t}\Delta t=-v(\alpha)\frac{d\alpha(t)}{dt}\Delta t=\mathrm{const.},$ (7)
where $v(\alpha)=\left.\frac{\partial\tilde{D}(\alpha,\alpha^{\prime})}{\partial\alpha^{\prime}}\right|_{\alpha^{\prime}=\alpha}$ represents the rate of change in the probability distributions of diffused data. Therefore, the noise-schedule function is determined to satisfy $-\frac{d\alpha(t)}{dt}\propto v(\alpha)^{-1}$. The proportionality coefficient can be calculated using the boundary conditions: $\alpha(0)=\alpha_{\max}$ and $\alpha(1)=\alpha_{\min}$. Finally, we obtain the noise-schedule function as follows:
@@
Here, we introduce a hyperparameter $\xi>0$ to control the dependence of $\alpha(t)$ on $v(\alpha)$. As $\xi$ is increased, more timesteps are allocated to the region where $v(\alpha)$ is large.
Similar to NCSN (score_matching1, ) inspired by simulated annealing (sa_original, ), our formulation draws inspiration from the adiabatic theorem (adiabatic_theorem, ) in quantum annealing (qa_original, ). Quantum annealing is an algorithm for solving combinatorial optimization problems, where the target cost function to be optimized is interpolated with an initial cost function with a trivial optimal solution. The initial state of the optimization is set to the optimal solution of the initial cost function. By tracing the optimal solutions of the instantaneous cost functions during the interpolation, the optimal solution of the target cost function is then obtained. In quantum annealing, accurately tracing the optimal solution of an instantaneous cost function is crucial, and the adiabatic theorem derives the traceability of the optimal solution at each step. Let $\alpha(t)$ be the weight of the target cost function and $1-\alpha(t)$ be the weight of the initial cost function. In this case, the inverse traceability can be expressed in the form of $\frac{d\alpha(t)}{dt}v(\alpha)$, which is the same form as Eq. (7). It has also been verified that an optimal functional form of $\alpha(t)$ is obtained by minimizing the maximum value of the inverse traceability grover_optimal , which ensures a consistent traceability throughout the interpolation process. In CRS, we regard the probability-distributional change as the inverse traceability and minimize the maximum distance for determining the noise schedule.
4.3 Distance Metrics
Any distance metric is applicable to CRS. We introduce three distance metrics.
Distance metric based on FID: FID is a widely used distance metric to evaluate the sample fidelity in the field of image generation. The FID calculates the Fréchet distance fid_score under a normality assumption in the vision-relevant feature space embedded using the Inception-V3 model inception_v3 , which is trained on the ImageNet classification. When FID is used as the distance metric for the pixel-space diffusion models, CRS allocates more timesteps to the region where the sample fidelity significantly deteriorates in the diffusion process ($\alpha\simeq 1$). A weakness of FID-based distance metric is that it cannot accurately capture the probability-distributional change in the region with high-noise levels ($\alpha\simeq 0$), because the feature model is not trained on diffused images.
It is important to note that, in the latent-space diffusion models, the probability-distributional change must be measured in the latent space embedded using the autoencoder (the approximation in Eq. (3) is applied to the diffusion process defined in the latent space). We trained ResNet50 resnet on ImageNet classification in the latent space embedded by the autoencoder and used it as the feature model for computing FID.
We refer to $v(\alpha)$ computed using FID as $v_{\mathrm{FID}}(\alpha)$. Once $v_{\mathrm{FID}}(\alpha)$ is computed, we can obtain the noise schedule using Eq. (8). For the detailed computation procedure of $v_{\mathrm{FID}}(\alpha)$, refer to Appendix C.1.
Distance metric based on data prediction: We introduce a distance metric that uses a trained diffusion model as follows:
$v_{x}(\alpha)=\sqrt{\left.\frac{\partial\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})}{\partial\alpha^{\prime}}\right|_{\alpha^{\prime}=\alpha}}\ \ \mathrm{where}\ \ \overline{D^{2}_{x}}(\alpha,\alpha^{\prime})=\mathbb{E}_{\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}}}\left[\left\|\bm{x}_{\theta}(\bm{x}_{\alpha},\alpha)-\bm{x}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})\right\|_{2}^{2}\right].\$ (9)
Here, $\bm{x}_{\alpha}$ and $\bm{x}_{\alpha^{\prime}}$ are sampled from $q(\bm{x}_{\alpha},\bm{x}_{\alpha}^{\prime})=q(\bm{x}_{\alpha}^{\prime}|\bm{x}_{\alpha})q(\bm{x}_{\alpha}|\bm{x}_{0})q(\bm{x}_{0})$, and $\bm{x}_{\theta}(\bm{x},\alpha)$ is the data prediction, which is related to the noise prediction $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$ as follows: $\bm{x}=\alpha\bm{x}_{\theta}(\bm{x},\alpha)+\sigma\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$. This distance metric aims at allocating more computational resources to the region where the data prediction changes rapidly. As shown in Appendix B, we can interpret this distance metric as a simplified version of the Kullback-Leibler divergence upper bound (KLUB) introduced in AYS. This distance metric can be used not only for determining the sampling schedule after training a diffusion model, but also for adaptively optimizing the training schedule during the training process. Pseudocodes for optimizing training and sampling schedules using $v_{x}(\alpha)$ are shown in Appendix C.2.
We can interpret that $v_{x}(\alpha)$ measures the probability-distributional change of diffused data. The data prediction $\bm{x}_{\theta}(\bm{x},\alpha)$ is given by the weighted sum of the data points $\bm{x}_{0}$ statistical_mechanics . The weight of each data point depends on the noise level $\alpha$. The weights of all data points are uniform when $\alpha=0$, while a small number of data points contribute to $\bm{x}_{\theta}(\bm{x},\alpha)$ when $\alpha\simeq 1$. The changes in these weights represent how the trajectories in the diffusion process gradually diverge as the noise level decreases. Therefore, we can interpret $v_{x}(\alpha)$ as measuring the probability-distributional change of diffused data.
Unlike the FID-based distance metric, $v_{x}(\alpha)$ can capture distributional changes even in regions with high-noise levels. However, $v_{x}(\alpha)$, which is related to the KLUB, is not necessarily optimal for improving sample fidelity at sampling. Therefore, as discussed in the next subsection, combining multiple distance metrics in a complementary manner enhances the performance of CRS.
Distance metric based on noise prediction: In the experiments presented in Sec. 5, we trained the noise prediction model. In this case, it is also considered effective to use noise predictions to measure the probability-distributional change as follows:
$v_{\varepsilon}(\alpha)=\sqrt{\left.\frac{\partial\overline{D^{2}_{\varepsilon}}(\alpha,\alpha^{\prime})}{\partial\alpha^{\prime}}\right|_{\alpha^{\prime}=\alpha}}\ \ \mathrm{where}\ \ \overline{D^{2}_{\varepsilon}}(\alpha,\alpha^{\prime})=\mathbb{E}_{\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}}}\left[\left\|\bm{\varepsilon}_{\theta}(\bm{x}_{\alpha},\alpha)-\bm{\varepsilon}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})\right\|_{2}^{2}\right].$ (10)
Hereafter, we refer to CRS using $v_{\mathrm{FID}}(\alpha)$, $v_{x}(\alpha)$, and $v_{\varepsilon}(\alpha)$ for optimizing noise schedules as CRS-$v_{\mathrm{FID}}$, CRS-$v_{x}$, and CRS-$v_{\varepsilon}$, respectively.
4.4 Combining Multiple Distance Metrics
As explained in the previous subsection, each distance metric has its own strengths an weaknesses. Therefore, it is considered effective to complementarily combining multiple distance metrics for optimizing noise schedules. We introduce a strategy for combining them as follows:
Here, $M$ is the number of distance metrics to be combined, $v_{m}(\alpha)$ is computed for each distance metric, $w_{m}$ is the weight for $v_{m}(\alpha)$ satisfying $\sum_{m=1}^{M}w_{m}=1$, and $\xi_{m}$ controls the dependence on $v_{m}(\alpha)$. Note that the value range of $v_{m}(\alpha)$ can significantly vary depending on the distance metric. Therefore, normalization of $v_{m}(\alpha)$ is required before combining them, as shown in Eq. (11).
Using this formulation, it is also possible to combine conventional noise schedules with CRS based noise schedules by converting $\alpha(t)$ to $v(\alpha)$. For example, when $\alpha(t)=\cos\left(\frac{\pi t}{2}\right)$, $v(\alpha)$ is obtained as follows:
$v(\alpha)\propto-\left(\frac{d\alpha(t)}{dt}\right)^{-1}=\frac{2}{\pi}\sin^{-1}\left(\frac{\pi t}{2}\right)=\frac{2}{\pi\sqrt{1-\alpha^{2}}}.$ (12)
This result suggests that every noise schedule can be interpreted as being optimized using CRS with its own distance metric, providing a unified perspective on the design of noise schedules.
5 Experiments
We experimentally demonstrated that CRS broadly improves the performance of diffusion models. Below, we provide a brief explanation of each component of the diffusion model.
Dataset: We use six image datasets: LSUN (Church, Bedroom, Horse) dataset_lsun , ImageNet dataset_imagenet , FFHQ dataset_ffhq , and CIFAR10 dataset_cifar10 . LSUN and FFHQ are used for unconditional image generation, and ImageNet and CIFAR10 for class-conditional image generation.
Noise Prediction Model: We use the U-Net architecture introduced in the ablated diffusion model (ADM) adm as the noise prediction model (see Table 8 in Appendix E for hyperparameter settings).
Training Schedule: We use linear ddpm_original , shifted cosine shifted_cosine , adaptive noise schedule proposed in VDM++ vdmpp , and CRS.
Loss Function: Except for the adaptive noise schedule of VDM++, we use the following simplified loss function introduced in DDPMs:
where $\mathcal{D}$ denotes all data points in the training dataset, $\alpha(t)$ is the training schedule, and $\bm{x}=\alpha(t)\bm{x}_{0}+\sqrt{1-\alpha(t)^{2}}\bm{\varepsilon}$.
When using the adaptive noise schedule of VDM++, the following loss function is applied:
$L=\frac{1}{2}\mathbb{E}_{\bm{x}_{0}\sim\mathcal{D},\lambda\sim p(\lambda),\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})}\left[\frac{\omega(\lambda)}{p(\lambda)}\left\|\bm{\varepsilon}_{\theta}\left(\bm{x},\sqrt{\mathrm{sigmoid}(\lambda)}\right)-\bm{\varepsilon}\right\|_{2}^{2}\right],$ (14)
where $\lambda=\log\left(\frac{\alpha^{2}}{\sigma^{2}}\right)$ is the log signal-to-noise ratio (log-SNR), $\omega(\lambda)$ is the weight for each noise level, and $p(\lambda)=-\frac{d\lambda}{dt}$ is the noise schedule for training. We use the EDM monotonic function as $\omega(\lambda)$, and $p(\lambda)$ is updated at each iteration to approximately satisfy the following condition:
Sampling Schedule: We use linear, shifted cosine, EDM sampling edm and CRS. In pixel-space diffusion models, we additionally evaluate CRS-$v_{x}+v_{\mathrm{FID}}$, which combines $v_{x}(\alpha)$ and $v_{\mathrm{FID}}(\alpha)$ for optimizing noise schedules. The reason for selecting the combination of $v_{x}(\alpha)$ and $v_{\mathrm{FID}}(\alpha)$ is that, as shown in Sec. 5.4, $v_{x}(\alpha)$ and $v_{\mathrm{FID}}(\alpha)$ have a complementary relationships. Recent studies edm ; vdmpp have shown that using different noise schedules for training and sampling can lead to improved performance. Therefore, in our experiments, we evaluate the performance of the diffusion models using arbitrary pairs of training and sampling schedules. Note that we can obtain different noise schedules for training and sampling using CRS, because it is capable of handling arbitrary distance metrics. Details of the conventional noise schedules used in our experiments can be found in Table 10 in Appendix E.
@@ 5 Experiments
Sampling Methods: Images are mainly generated using two stochastic samplers and three deterministic samplers. For the stochastic samplers, we use Stochastic DDIM ($\eta=1$) ddim_original and SDE-DPM-Solver++(2M) sampler_dpmpp . For the deterministic samplers, we use DDIM ddim_original , DPM-Solver++(2M) sampler_dpmpp , and pseudo numerical methods for diffusion models (PNDMs) sampler_pndm . At small NFEs, we additionally use improved PNDM (iPNDM) sampler_ipndm and a unified predictor-corrector framework (UniPC) sampler_unipc .
@@
Evaluation Metrics: We compute four metrics: FID fid_score , spatial FID (sFID) spatial_fid_score , and improved precision and recall recall_precision1 ; recall_precision2 . FID, sFID, and precision measure sample fidelity, while recall evaluates mode coverage.

### v4

@@ 4.2 Constant Rate Scheduling
$D(t,t+\Delta t)\simeq\left.\frac{\partial D(t,t^{\prime})}{\partial t^{\prime}}\right|_{t^{\prime}=t}\Delta t=v(\alpha)\frac{d\alpha(t)}{dt}\Delta t=\mathrm{const.},$ (8)
where $v(\alpha)$ represents the rate of change in the probability distribution of the diffused data.
Therefore, the noise schedule must satisfy $\frac{d\alpha(t)}{dt}\propto v(\alpha)^{-1}$. The proportionality constant is computed from the boundary conditions $\alpha(0)=\alpha_{\max}$ and $\alpha(1)=\alpha_{\min}$, yielding:
$\frac{d\alpha(t)}{dt}=Cv(\alpha)^{-\xi},\quad\mathrm{where}\quad C=\int_{\alpha_{\min}}^{\alpha_{\max}}v(\alpha)^{\xi}d\alpha.$ (9)
Here, we introduce a hyperparameter $\xi>0$ to control the dependence of $\alpha(t)$ on $v(\alpha)$. Larger values of $\xi$ allocate more timesteps to regions where $v(\alpha)$ is large, i.e., where the distribution changes rapidly.
The noise schedule optimization using CRS consists of the following steps:
  1. 1.
    Choose a discrepancy measure to quantify distributional change.
  2. 2.
    Compute $v(\alpha)$ using the selected discrepancy measure.
  3. 3.
    Solve Eq. (9) to obtain $\alpha(t)$.
  4. 4.
    Apply the resulting schedule as follows:
    • For training, use $\alpha(t)$ directly.
    • For sampling, discretize the schedule uniformly in time: $\{\alpha(t/T)|t=0,1,...,T\}$ where $T$ is the number of sampling steps.
We propose three discrepancy measures for computing $v(\alpha)$, which are detailed in Section 4.3.
The derivation of Eqs. (7) and (8) relies on a first-order Taylor expansion with respect to $\Delta t$, which is accurate when the NFE is sufficiently large. However, when optimizing the sampling schedule at small NFEs, $\Delta t$ becomes large and the approximation error may no longer be negligible. Empirically, we observe that CRS with $\xi=1$ remains effective for moderate NFEs, while the importance of tuning $\xi$ increases as the NFE decreases (see Table 27 in Appendix F.2). We attribute this increased sensitivity to the growing discretization error of the first-order expansion.
For small NFEs, the optimal value of $\xi$ tends to be greater than $1$, resulting in more timesteps being allocated to regions where $v(\alpha)$ is large. Since discretization error is expected to be more pronounced in regions with larger $v(\alpha)$, it is intuitively reasonable that values of $\xi>1$ become optimal when the NFE is small.
Although introducing $\xi$ partially compensates for the discretization error, it does not fully resolve the mismatch between the continuous-time formulation and the discrete-time implementation. A fully principled solution would require directly optimizing the sampling schedule in the discrete-time domain, without relying on the continuous-time assumption underlying Eq. (7). We consider this an important direction for future work.
Similar to NCSNs inspired by simulated annealing (Kirkpatrick et al., 1983), our formulation draws additional motivation from the adiabatic theorem (Morita and Nishimori, 2008) in quantum annealing (Kadowaki and Nishimori, 1998). Quantum annealing solves combinatorial optimization problems by interpolating between a target cost function and an initial one that has a trivial optimum. The optimization starts from the known optimum of the initial function and aims to trace the instantaneous optimums along the interpolation path. It is crucial to accurately trace the optimum of the instantaneous cost function, and the adiabatic theorem provides the traceability of the optimum at each step. Let $\alpha(t)$ be the weight of the target cost function and $1-\alpha(t)$ be the weight of the initial cost function. In this setting, the inverse traceability is expressed in the form of $\frac{d\alpha(t)}{dt}v(\alpha)$, which takes the same form as Eq. (8). Prior work has shown that minimizing the maximum value of this inverse traceability yields the optimal $\alpha(t)$ (Roland and Cerf, 2002), that ensures consistent traceability throughout the interpolation. In CRS, we regard the rate of probability-distributional change as the inverse traceability and minimize its maximum value to derive an optimal noise schedule.
4.3 Discrepancy Measures
Any discrepancy measure is applicable to CRS. We introduce three discrepancy measures.
4.3.1 Discrepancy measure based on FID
FID is a widely used metric for evaluating sample fidelity in the field of image generation. It computes the Fréchet distance (Heusel et al., 2017) between the feature distributions of real and generated images under a Gaussian assumption, where the features are extracted using the Inception-V3 network (Szegedy et al., 2016) trained on ImageNet classification.
When used as a discrepancy measure in our framework, $v(\alpha)$ tends to take large values in the region where sample fidelity significantly degrades (i.e., $\alpha\simeq 1$). As a result, CRS allocates more timesteps to this region, thereby improving sample fidelity.
However, a limitation of the FID-based measure is that it fails to accurately capture distributional changes in high-noise regions ($\alpha\simeq 0$), since the feature extractor is not trained on heavily diffused images and may not provide meaningful representations in that regime. Therefore, while the FID-based measure is effective for capturing fidelity-related changes in low-noise regions, it may benefit from being complemented by other discrepancy measures that can better capture high-noise dynamics.
Once the forward process is defined, $v(\alpha)$ can be computed by simulating the forward diffusion using the training dataset (see Algorithm 1 in Appendix C.1). Since this procedure is independent of model parameters, it only needs to be executed once prior to training the diffusion model.
The computation time of $v(\alpha)$ is proportional to the number of training images, which may be prohibitively expensive for large-scale training datasets. To address the scalability issue for large-scale datasets, we introduce a practical alternative in Sec. 5.7.
We refer to $v(\alpha)$ computed using FID as $v_{\mathrm{FID}}(\alpha)$.
4.3.2 Discrepancy measure based on data prediction
We introduce a discrepancy measure that leverages a trained diffusion model as follows:
$v_{x}(\alpha)=\sqrt{\left.\frac{\partial\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})}{\partial\alpha^{\prime}}\right|_{\alpha^{\prime}=\alpha}},\quad\mathrm{where}\quad\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})=\mathbb{E}_{\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}}\sim q(\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}})}\left[\left\|\bm{x}_{\theta}(\bm{x}_{\alpha},\alpha)-\bm{x}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})\right\|_{2}^{2}\right].$ (10)
Here, $q(\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}})$ is the joint distribution of diffused data at different noise levels, and $\bm{x}_{\theta}(\bm{x},\alpha)$ denotes the data prediction, which is related to the noise prediction $\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$ via $\bm{x}=\alpha\bm{x}_{\theta}(\bm{x},\alpha)+\sigma\bm{\varepsilon}_{\theta}(\bm{x},\alpha)$.
This measure is designed to allocate more computational resources to regions where the data prediction changes rapidly with respect to $\alpha$. As shown in Appendix B, $v_{x}(\alpha)$ can be interpreted as a simplified variant of KLUB, introduced in AYS. Importantly, $v_{x}(\alpha)$ can be used not only to determine the sampling schedule after training, but also to adaptively optimize the training schedule during learning.
We further interpret $v_{x}(\alpha)$ as measuring the probability-distributional change of the diffused data. According to prior work (Ambrogioni, 2024), the data prediction $\bm{x}_{\theta}(\bm{x},\alpha)$ can be viewed as a weighted average over training data points $\bm{x}_{0}$, where the weighting depends on the noise level $\alpha$. When $\alpha=0$, the weights are approximately uniform across the dataset. As $\alpha\rightarrow 1$, only a few data points contribute significantly, indicating sharper and more specific representations. The change in these implicit weights reflects how sample trajectories diverge as noise decreases, and thus $v_{x}(\alpha)$ provides a proxy for the rate of distributional change.
The pseudocode for optimizing training and sampling schedules is provided in Appendix C.2. For sampling schedule optimization, the computation time of $v_{x}(\alpha)$ is negligible compared to the overall training time. In contrast, training schedule optimization requires updating $\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})$ during training as the model parameters evolve, which increases the total training time by approximately 20–30%. However, our experiments show that the schedule typically stabilizes early in training. Therefore, fixing the schedule after a few epochs or reducing the update frequency of $\overline{D^{2}_{x}}(\alpha,\alpha^{\prime})$ may mitigate the computational cost without compromising performance. Developing a more efficient implementation remains an important direction for future work.
4.3.3 Discrepancy measure based on noise prediction
In the experiments presented in Sec. 5, we trained the noise prediction model. In this setting, it is also effective to use the noise predictions themselves to quantify distributional change, leading to the following definition:
$v_{\varepsilon}(\alpha)=\sqrt{\left.\frac{\partial\overline{D^{2}_{\varepsilon}}(\alpha,\alpha^{\prime})}{\partial\alpha^{\prime}}\right|_{\alpha^{\prime}=\alpha}},\quad\mathrm{where}\quad\overline{D^{2}_{\varepsilon}}(\alpha,\alpha^{\prime})=\mathbb{E}_{\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}}\sim q(\bm{x}_{\alpha},\bm{x}_{\alpha^{\prime}})}\left[\left\|\bm{\varepsilon}_{\theta}(\bm{x}_{\alpha},\alpha)-\bm{\varepsilon}_{\theta}(\bm{x}_{\alpha^{\prime}},\alpha^{\prime})\right\|_{2}^{2}\right].$ (11)
We refer to CRS using $v_{\mathrm{FID}}(\alpha)$, $v_{x}(\alpha)$, and $v_{\varepsilon}(\alpha)$ as $v_{\mathrm{FID}}, CRS$-$v_{x}$, and CRS-$v_{\varepsilon}$, respectively, depending on the choice of discrepancy measure used to compute $v(\alpha)$.
4.4 Combining Multiple Discrepancy Measures
Each discrepancy measure has its own strengths and weaknesses. For example, CRS-$v_{\mathrm{FID}}$ effectively captures distributional changes relevant to sample fidelity, but fails to represent changes accurately in high-noise regions ($\alpha\simeq 0$), as the feature extractor is not trained on diffused images. On the other hand, CRS-$v_{x}$ captures distributional change even in high-noise regions, but it is not necessarily optimal for improving sample fidelity. Therefore, it is considered effective to complementarily combining multiple discrepancy measures for optimizing noise schedules.
We introduce the following strategy for combining multiple discrepancy measures:
where $M$ is the number of discrepancy measures, $v_{m}(\alpha)$ is the rate of change computed using the $m$-th measure, $w_{m}$ is its corresponding weight (satisfying $\sum_{m=1}^{M}w_{m}=1$), and $\xi_{m}$ controls the emphasis on $v_{m}(\alpha)$. Since the value ranges of different discrepancy measures can vary significantly, we normalize each $v_{m}(\alpha)$ as in Eq. (12) before combining them.
This formulation also enables conventional noise schedules to be interpreted and incorporated within the CRS framework. By converting a given noise schedule $\alpha(t)$ into its corresponding form of $v(\alpha)$, we can treat it as an implicit discrepancy measure. For example, if $\alpha(t)=\cos\left(\frac{\pi t}{2}\right)$, then:
$v(\alpha)\propto-\left(\frac{d\alpha(t)}{dt}\right)^{-1}\propto\sin^{-1}\left(\frac{\pi t}{2}\right)=\frac{1}{\sqrt{1-\alpha^{2}}}.$ (13)
This observation implies that any noise schedule can be viewed as the optimization result of CRS with an associated (possibly implicit) discrepancy measure, thus providing a unified perspective on the design of noise schedules.
While combining multiple discrepancy measures increases the flexibility of CRS, it also introduces additional hyperparameters. In our experiments, we found that tuning these hyperparameters is required only when optimizing sampling schedules in pixel-space diffusion models. In all other cases—training schedules for both pixel-space and latent-space models, and sampling schedules for latent-space models—CRS-$v_{x}$ with $\xi=1$ consistently performed well and can serve as a default setting.
In the next section, we show that CRS-$v_{x}+v_{\mathrm{FID}}$, which combines $v_{x}(\alpha)$ and $v_{\mathrm{FID}}(\alpha)$, is effective for optimizing the sampling schedule of pixel-space diffusion models. CRS-$v_{x}+v_{\mathrm{FID}}$ has four hyperparameters: $w_{x}$, $w_{\mathrm{FID}}$, $\xi_{x}$, and $\xi_{\mathrm{FID}}$. The procedure for tuning these hyperparameters is described in Appendix D, and the hyperparameter settings used in all experiments are summarized in Table 12. For configurations with $\mathrm{NFE}\leq 50$, Table 12 shows that $w_{x}=w_{\mathrm{FID}}=0.5$ is optimal in most cases. These observations suggest that, in practical use, it is generally sufficient to tune only $\xi_{x}$ and $\xi_{\mathrm{FID}}$.
5 Experiments
We experimentally demonstrate that CRS broadly improves the performance of diffusion models.
5.1 Experimental Settings
Below, we provide a brief overview of the components and settings used in our experiments.
Dataset: We use six image datasets: LSUN (Church, Bedroom, Horse) (Yu et al., 2016), ImageNet (Deng et al., 2009), FFHQ (Karras et al., 2019), and CIFAR10 (Krizhevsky, 2009). LSUN, FFHQ, and CIFAR10 are used for unconditional image generation, while ImageNet is used for class-conditional image generation.
Noise Prediction Model: We adopt the U-Net architecture introduced in the ablated diffusion model (ADM) (Dhariwal and Nichol, 2021) as the noise prediction model. Hyperparameter settings are listed in Table 13 (Appendix E).
Training Schedule: We compare four training noise schedules: linear (Ho et al., 2020), shifted cosine (Hoogeboom et al., 2023), the adaptive schedule proposed in VDM++ (Kingma and Gao, 2023), and our proposed CRS. The CRS hyperparameter $\xi$ is fixed to $1$ in all experiments.
Loss Function: Except for the adaptive schedule in VDM++, we use the simplified loss function from DDPMs:
where $\bm{x}=\alpha(t)\bm{x}_{0}+\sqrt{1-\alpha(t)^{2}}\bm{\varepsilon}$, $\mathcal{D}$ denotes the training dataset, and $\alpha(t)$ is the training schedule.
When using the adaptive schedule from VDM++, the following loss function is used:
$L=\frac{1}{2}\mathbb{E}_{\bm{x}_{0}\sim\mathcal{D},\lambda\sim p(\lambda),\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I})}\left[\frac{\omega(\lambda)}{p(\lambda)}\left\|\bm{\varepsilon}_{\theta}\left(\bm{x},\sqrt{\mathrm{sigmoid}(\lambda)}\right)-\bm{\varepsilon}\right\|_{2}^{2}\right],\\$ (15)
where $\lambda=\log\left(\frac{\alpha^{2}}{\sigma^{2}}\right)$ is the log signal-to-noise ratio (log-SNR), $\omega(\lambda)$ is the weighting function, and $p(\lambda)=-\frac{d\lambda}{dt}$ denotes the noise schedule. We adopt the EDM monotonic function as $\omega(\lambda)$, and update $p(\lambda)$ at each iteration to approximately satisfy:
@@ 5.1 Experimental Settings
Sampling Methods: To assess the generality across samplers, we employ two stochastic samplers and three deterministic samplers. The stochastic samplers include Stochastic DDIM ($\eta=1$) (Song et al., 2021a) and SDE-DPM-Solver++(2M) (Lu et al., 2022b). The deterministic samplers include DDIM (Song et al., 2021a), DPM-Solver++(2M) (Lu et al., 2022b), and pseudo numerical methods (PNDM) (Liu et al., 2022). At low NFE regimes, we also evaluate improved PNDM (iPNDM) (Zhang and Chen, 2023) and the unified predictor-corrector framework (UniPC) (Zhao et al., 2023).
@@
Evaluation Metrics: We use four evaluation metrics: FID (Heusel et al., 2017), spatial FID (sFID) (Nash et al., 2021), and improved precision and recall (Sajjadi et al., 2018; Kynkäänniemi et al., 2019). FID, sFID, and precision measure sample fidelity, while recall quantifies mode coverage.